OpenStudy (anonymous):

Graph the following equation. Begin by completing the square on the first two terms: y=x^2+8x+5

OpenStudy (anonymous):

\[x^2+8x+5=x^2+8x+16-9=(x+4)^2-9\] Identify the vertex point, and you're pretty much set to graph.

OpenStudy (anonymous):

How do I find the points? I'm trying to finish the quiz before I do the homework, so I have no idea how to do this, sorry lol

OpenStudy (anonymous):

Heeeey, @Jack1 !! :)

OpenStudy (jack1):

haya cookie ;)

OpenStudy (anonymous):

I need super help tonight, I'm trying to finish like 3 homeworks and a quiz tonight. I'm doing the quiz first :/

OpenStudy (jack1):

for this, i still do it the long way, solve your x intercepts, find ur turning point, and then pick a 3rd y value based on that, then connect the 5 points and ur there, im sure theres an easier way tho, probably to do with vertex point,

OpenStudy (anonymous):

Sounds difficult lol I haven't done the homework for this chapter so I have no idea what to do

OpenStudy (jack1):

wow, gud luck, im just takin a 5 min breather, i have to finish another assignment b4 wednesday

OpenStudy (jack1):

ok, so have u done derivatives b4?

OpenStudy (anonymous):

I have to finish all of my work by Tuesday! Finals are Wednesday >.<

OpenStudy (anonymous):

uhhh, probably

OpenStudy (jack1):

damn... well, hang on... googling good website to explain better

OpenStudy (jack1):

sweet, got it: http://www.mathwarehouse.com/geometry/parabola/standard-and-vertex-form.php go here and look at the pics on this website

OpenStudy (jack1):

what @SithsAndGiggles was helping with before was putting the equation in vertex form, as that will easily give u the x and y coordinate of ur turning point

OpenStudy (anonymous):

So I graph 4 -9?

OpenStudy (anonymous):

Or something like that...

OpenStudy (jack1):

k sorry g2g cookie as i doan have much time but think it will look something like this: hmmm... wait: think @SithsAndGiggles made a little typo above , should proly be -11, not -9

OpenStudy (jack1):

\[\Large x^2+8x+5=x^2+8x+16-11=(x+4)^2-11\] 16 - 9 = 7 16 - 11 = 5 our original equation is for 5

OpenStudy (jack1):

so in vertex form: \[\large y= a(x-h) +k\]\[ \large y =(x+4)^2-11\] so (-4) = h and (- 11) = k so that's ur turning point use quadratic formula to solve for x intercepts then pick a point like y = 10 and solve for the 2 x points now connect the dots and ur done :)

OpenStudy (jack1):